# Leveraging Multi-facet Paths for Heterogeneous Graph Representation Learning

### v2

@@ Abstract.
Recent advancements in heterogeneous GNNs have enabled significant progress in embedding nodes and learning relationships across diverse tasks. However, traditional methods rely heavily on meta-paths grounded in node types, which often fail to encapsulate the full complexity of node interactions, leading to inconsistent performance and elevated computational demands. To address these challenges, we introduce MF2Vec, a novel framework that shifts focus from rigid node-type dependencies to dynamically exploring shared facets across nodes, regardless of type. MF2Vec constructs multi-faceted paths and forms homogeneous networks to learn node embeddings more effectively. Through extensive experiments, we demonstrate that MF2Vec achieves superior performance in node classification, link prediction, and node clustering tasks, surpassing existing baselines. Furthermore, it exhibits reduced performance variability due to meta-path dependencies and achieves faster training convergence. These results highlight its capability to analyze complex networks comprehensively. The implementation of MF2Vec is publicly available at https://anonymous.4open.science/r/MF2Vec-6ABC.
@@ 1. Introduction
Figure 2 demonstrates that even when following the same meta-path, nodes can be grouped based on different facets. For example, the paths “New York-L. Dicaprio-The Wolf of Wall Street” and “Interstellar-Anne H.-Intern” may simply be grouped as “Movie-Actor-Movie” under meta-path-based methods. However, with a multi-facet approach, the path “New York-L. Dicaprio-The Wolf of Wall Street” can be considered from the perspective of filming location, while the path “Interstellar-Anne H.-Intern” may reflect similarities in IMDb ratings. This approach allows the model to comprehensively capture latent and rich information within the graph, effectively overcoming the limitations of traditional meta-path-based methods.
In MF2Vec, paths are extracted from the adjacency matrix of a heterogeneous graph through random walks without distinguishing them by predefined meta-path types. The node embeddings of the extracted paths are projected into a shared facet space of dimensions. Gumbel Softmax and an attention mechanism are employed to select the most relevant facets for intermediate nodes, ensuring a discrete yet differentiable selection process. These facets characterize the relationships between endpoint nodes and compute path weights that reflect their importance for node connections. Finally, a homogeneous network is constructed by integrating same-typed nodes and fed into a Graph Convolutional Network (GCN) (Kipf and Welling, 2016), which aggregates multi-facet vectors along the paths to generate latent node representations. Key contributions of this work are summarized as follows:
@@
Efficiency and Performance Enhancement: Reduction in computational costs and enhanced performance for node classification, node clustering, and link prediction tasks
@@
2. Related Work
2.1. Heterogeneous Graph Neural Network (HGNN)
2.1.1. Relation-based HGNNs
2.1.2. Metapath-based HGNNs
2.2. Multiple-Vector Network Embedding
@@ 3. Preliminary
This section provides definitions of terminologies used in the current study. The notations and their explanations are summarized in Table 1.
Definition 1 (Heterogeneous Graph). A heterogeneous graph is characterized by a node type mapping function and an edge type mapping function , where and are sets of node and edge types, respectively, with .
@@
Definition 3 (Metapath Schema). A metapath schema is a sequence of node types connected by edge types . This schema represents the structural pattern of a metapath and illustrates how different node types are related through specific types of relationships defined by the nodes and edges.
@@ 4. Methodology
In this section, we describe our approach, as shown in Figure 3. We begin by extracting paths between homogeneous end nodes using random walks, similar to traditional metapath methods (Dong et al., 2017; Fu et al., 2017). We then project the features of intermediate nodes into multiple facets and aggregate these facets to create multi-facet embedding representations. These representations are used as edge features between path-guided neighbors and the target node. The model pseudocode is provided in Algorithm 1, and the notations are summarized in Table 1.
@@ 4.1. Extracting Paths Using Random Walk
Given a node type , our method constructs a homogeneous subgraph that identifies direct connections between nodes of type , the same as traditional metapath-based models. Instead of relying on metapath schemas such as APA or APCPA, the method generates a path starting from a node of type and applies a random walk, traversing randomly selected adjacent nodes until encountering a type A node. This approach captures relationships within a specified path length limit, effectively mapping connections within the nodes of a type without being constrained by predefined metapath patterns. Details of the Random Walk algorithm are provided in Algorithm 2.
@@ 4.2. Transforming Single Embeddings into Multi-Facet
Our model captures inherent information from all node types before generating a facet embedding. Initially, node information is exchanged through a GNN warm-up phase, which facilitates the aggregation of neighborhood information and enriches the initial node embeddings with structural and semantic context. The node embedding, , is defined during this warm-up phase and serves as the foundation for training the main model. The resulting node embedding is then refined into a multi-dimensional embedding to capture diverse information, as detailed in Equation 1.
Here, represents the multi-facet embedding of node within the -th facet, where is the corresponding weight matrix with dimensions for the -th facet, and represents the total number of facets. The individual facet embeddings are subsequently concatenated into a unified multi-facet embedding, , providing a comprehensive representation.
@@ 4.4. Discriminative Facet Consideration using Gumbel Softmax
We utilize the Gumbel Softmax (Jang et al., 2016) function to introduce randomness and uncertainty into facet selection, which helps the model explore diverse facet influences in complex relationships. The temperature parameter allows us to control the balance between exploration and exploitation. This approach enables us to capture nuanced facet interactions and dependencies, making our model more adaptable and robust for multi-facet data. Even if the random walk algorithm extracts the paths that are little related to the end nodes, we mitigate their impact by assigning higher weights to the most crucial components using Gumbel softmax. As a result, we effectively operate within constrained environments, even without extracting meaningful paths.
In the given equation, represents Gumbel Softmax. is the final embedding for that combines these probabilities.and is a matrix with dimensions .
@@ 4.5. Subgraph Construction
Our approach, however, avoids the need for multiple metapaths by leveraging multi-facet information from a single path between nodes. This results in a subgraph that provides richer information than traditional methods. The MF2Vec model efficiently learns complex node relationships by constructing subgraphs as described in Algorithm 3 incorporating multi-dimensional edge features which are multi-facet embedding for a more comprehensive node representation.
@@ 4.6. Message Passing
To extract node embeddings, we utilize GCN. In this process, we consider interactions between nodes and utilize path facets as edge features. The formula for calculating graph node embeddings is as follows:
@@ 4.7. Training Losses
We train our model for two tasks: node classification and link prediction. For node classification, the cross-entropy loss is computed over labeled nodes , measuring the disparity between true labels and predictions :
For link prediction, where node labels are unavailable, we minimize the contrastive loss using positive and negative node pairs:
Here, is the sigmoid function, ensuring effective learning for link prediction.
@@ 5. Experiment
In this part, we assess the performance and practicality of our introduced method through tests on link prediction, node classification, and node clustering tasks across five diverse graph datasets. The primary goals of these experiments are to address the following research questions.
RQ1. How well does MF2Vec perform on downstream tasks?
RQ2. How stable is the performance of MF2Vec in downstream tasks?
RQ3. How time-efficient is MF2Vec?
5.1. Experimental Settings
5.1.1. Datasets
We evaluate our model on six heterogeneous graph datasets: DBLP, ACM, IMDB, and Freebase for node classification and clustering, and Yelp and MovieLens for link prediction. Experiments include 1000 trials per node with a walk length of 5, covering typical metapaths (e.g., APA, APCPA) using consistent data splits. Dataset details are summarized in Table 2.
@@ 5.1.1. Datasets
5.1.2. Baselines
We compare MF2Vec against state-of-the-art heterogeneous graph neural networks (HGNNs), categorized as follows: (a) Relation-Based HGNNs, including RGCN, RGAT, HGT, HetSANN, and Simple-HGN; and (b) Metapath-Based HGNNs, comprising MP2Vec, HAN, MAGNN, HeCo, MHNF, HGMAE HMSG, and MECCH.
5.1.3. Graph Construction Settings
Following Algorithm 2, we ensured that the node types matched the target node and that the path length was at most 5. For each target node, we performed up to 1000 trials. The number of negative samples was set to 1.
5.1.4. Experimental Settings
To ensure a fair comparison among all heterogeneous graph neural networks (HGNNs), the node embedding dimension was fixed to 64. The learning rate was set to , weight decay to , and the dropout rate to 0.5. For each baseline model, hyperparameters such as batch size, attention heads, dropout rate, learning rate, and epochs—excluding the node embedding dimension—followed the default settings specified in the original papers. Training was conducted using the Adam optimizer with early stopping (patience of 20 epochs) to prevent overfitting. Specifically, for MF2Vec, the temperature () was set to 0.5, and the number of facets () was set to 5. To demonstrate generalization, the experiments were conducted using five seed values (1, 10, 100, 1000, and 10000), with the data split into train (80%), validation (10%), and test (10%). The model seeds were fixed during the experiments, and the average and standard deviation of the results were calculated for all seeds.
5.2. Downstream Task (RQ1)
@@
MF2Vec | 0.9910.002 | 0.9460.011 | 0.9430.010 | 0.9540.008 | 0.8430.008 | 0.8370.008 | 0.7680.033 | 0.6030.041 | 0.5810.052 | 0.8280.023 | 0.6940.021 | 0.6290.047 |
5.2.1. Node Classification
In our node classification experiments, we convert nodes into low-dimensional vector embeddings and apply a Softmax function to obtain probability distributions over label categories. We evaluate the model using Micro-F1, Macro-F1, and AUC metrics, providing a balanced assessment of classification performance. These metrics are used to compare our model, MF2Vec, against baseline models on the DBLP, ACM, IMDB and Freebase datasets, with results presented in Table LABEL:table_1.
Our approach consistently outperforms baseline models across all metrics, demonstrating the effectiveness of MF2Vec. The significant performance improvement, particularly on the IMDB dataset, highlights the model’s ability to capture complex relational data through multi-facet vectors. This suggests that MF2Vec effectively reveals nuanced relationships within heterogeneous information networks (HINs), enhancing understanding beyond node types.
5.2.2. Link Prediction
@@
In Movielens and Yelp, link prediction is framed as a binary classification task to determine the presence of an edge in the original graph. During training, validation, and testing, negative edges are created by replacing the destination node in positive edges with a randomly chosen node that is not connected to the source node. The link probability is computed using the dot product of node embeddings: , where is the sigmoid function.
Table 4 presents the AUC, Micro-F1, and Macro-F1 scores for our model and baselines on Yelp and Movielens. MF2Vec performs competitively on Yelp (against MECCH) and outperforms all baselines on MovieLens. These results highlight the advantage of using multi-facet vectors derived from node paths for link prediction, and demonstrate the efficacy of our model in updating embeddings without relying on node types.
@@
5.2.3. Node Clustering
5.3. Stability of performance (RQ2)
@@
DBLP | AUC | 0.144 | 0.201 | 0.118 | 0.064 |
Mi-F1 | 0.233 | 0.318 | 0.195 | 0.138 | |
Ma-F1 | 0.251 | 0.277 | 0.197 | 0.139 | |
ACM | AUC | 0.147 | 0.188 | 0.211 | 0.141 |
Mi-F1 | 0.083 | 0.106 | 0.076 | 0.070 | |
Ma-F1 | 0.102 | 0.088 | 0.066 | 0.051 | |
IMDB | AUC | 0.042 | 0.041 | 0.043 | 0.031 |
Mi-F1 | 0.029 | 0.049 | 0.067 | 0.028 | |
Ma-F1 | 0.032 | 0.030 | 0.032 | 0.020 |
This section emphasizes the stability advantage of leveraging multi-facets from paths in our model. Unlike metapath-based models that depend on predefined schemas (e.g., APA, APCPA), MF2Vec achieves consistent performance across different schema types, effectively capturing fine-grained information regardless of metapath type. To validate this, we compare single-metapath performance across all models, including MF2Vec, on the DBLP, ACM, and IMDB datasets.
Table 6 presents the standard deviations in node classification performance using a single metapath. While other models show significant performance variation depending on the chosen metapath, MF2Vec demonstrates remarkable stability. For the DBLP dataset, performance variations highlight the reliance of metapath-based models on predefined paths and the associated time and cost of analysis, which MF2Vec effectively mitigates.
5.4. Time Complexity (RQ3)
@@
The proposed model’s time complexity is , where is the number of nodes, is the average neighbors, is the number of facets, and is the embedding dimension. Attention adds , and graph convolution contributes , resulting in a total complexity of .
In Table 7, models like MECCH (Fu and King, 2024), HAN (Wang et al., 2019), and MAGNN (Fu et al., 2020) rely on metapaths, incurring higher costs due to growth with metapath length (). MF2Vec avoids this by leveraging facets (), leading to more efficient computation, satisfying .Additionally, we compare MF2Vec with relation-based methods like HetSANN (Hong et al., 2020) and SimpleHGN (Lv et al., 2021), which also avoid metapaths by focusing on edge types.
Figure 5 compares training times (log scale) and Macro F1 scores for MF2Vec and other models on DBLP, IMDB, and Movielens. MAGNN (Fu et al., 2020) and MP2Vec (Dong et al., 2017) are excluded from Movielens due to their excessive training times. MF2Vec achieves the best performance on node classification and link prediction tasks, with significantly lower training times. This efficiency arises from its dynamic use of path characteristics, avoiding the high costs of aggregating multiple paths as seen in MAGNN and MP2Vec, ensuring better scalability and faster convergence without compromising performance.

### v3

@@ Abstract.
Recent advancements in heterogeneous GNNs have enabled significant progress in embedding nodes and learning relationships across diverse tasks. However, traditional methods rely heavily on meta-paths grounded in node types, which often fail to encapsulate the full complexity of node interactions, leading to inconsistent performance and elevated computational demands. To address these challenges, we introduce MF2Vec, a novel framework that shifts focus from rigid node-type dependencies to dynamically exploring shared facets across nodes, regardless of type. MF2Vec constructs multi-faceted paths and forms homogeneous networks to learn node embeddings more effectively. Through extensive experiments, we demonstrate that MF2Vec achieves superior performance in node classification, link prediction, and node clustering tasks, surpassing existing baselines. Furthermore, it exhibits reduced performance variability due to meta-path dependencies and achieves faster training convergence. These results highlight its capability to analyze complex networks comprehensively. The implementation of MF2Vec is publicly available at https://github.com/kimjongwoo-cell/MF2Vec.
@@ 1. Introduction
Plot showing Macro-F1 and time complexity comparison between MECCH and MAGNN on DBLP dataset.
@@
Schema comparison on IMDB data: (A) predefined node-type schemas vs. (B) dynamic multi-facet path assignment.
Figure 2 shows that when following the same meta-path, nodes can be grouped based on different facets. For example, the paths “New York–L. Dicaprio–The Wolf of Wall Street” and “Interstellar–Anne H.–Intern” may simply be grouped as “Movie–Actor–Movie” under meta-path-based methods. However, with a multi-facet approach, the path “New York–L. Dicaprio–The Wolf of Wall Street” can be considered from the perspective of filming location, while the path “Interstellar–Anne H.–Intern” may reflect similarities in IMDb ratings. This approach allows the model to comprehensively capture latent and rich information within the graph, effectively overcoming the limitations of traditional meta-path-based methods.
In MF2Vec, paths are extracted from the adjacency matrix of a heterogeneous graph through random walks without distinguishing them by predefined meta-path types. The node embeddings of the extracted paths are projected into a shared facet space of dimensions. Gumbel Softmax and an attention mechanism are employed to select the most relevant facets for intermediate nodes, ensuring a discrete yet differentiable selection process. These facets characterize the relationships between endpoint nodes and compute path weights that reflect their importance for node connections. Finally, a homogeneous network is constructed by integrating same-typed nodes and fed into a GNN, which aggregates multi-facet vectors along the paths to generate latent node representations.
Key contributions of this work are summarized as follows:
@@
Efficiency and Performance Enhancement: Reduced computational costs and improved performance for node classification, node clustering, and link prediction tasks.
@@ 2. Preliminary
Definition 1 (Heterogeneous Graph). A heterogeneous graph is defined by a node type mapping function and an edge type mapping function , where and are sets of node and edge types, respectively, with .
@@
Definition 3 (Metapath Schema). A metapath schema is a sequence of node types connected by edge types . This represents the structural pattern of a metapath and illustrates how different node types are related through specific types of relationships defined by the nodes and edges.
@@ 3. Methodology
In this section, we describe our approach, as shown in Figure 3. We begin by extracting paths between homogeneous end nodes using random walks, similar to traditional metapath methods (Dong et al., 2017; Fu et al., 2017). We then project the features of intermediate nodes into multiple facets and aggregate these facets to create multi-facet embedding representations. These representations are used as edge features between path-guided neighbors and the target node. The model pseudocode is provided in Algorithm 1, and the notation is summarized in Table 1.
@@ 3.1. Extracting Paths Using Random Walk
Given a node type , our method constructs a homogeneous subgraph that identifies direct connections between nodes of type , the same as metapath-based models. Instead of relying on metapath schemas such as APA or APCPA, the method generates a path starting from a node of type and applies a random walk, traversing randomly selected adjacent nodes until encountering a type A node. This approach captures relationships within a specified path length limit, effectively mapping connections within the nodes of a type without being constrained by predefined metapath patterns. Details of the Random Walk algorithm are provided in Algorithm 2.
@@ 3.2. Transforming Single Embeddings into Multi-Facet
Our model captures inherent information from all node types before generating a facet embedding. Node information is exchanged through a GNN warm-up phase, which facilitates the aggregation of neighborhood information and enriches the initial node embeddings with structural and semantic context. The node embedding, , is defined during this warm-up phase and serves as the foundation for training the main model. The resulting node embedding is then refined into a multi-dimensional embedding to capture diverse information, as detailed in Equation 1.
Here, denotes the embedding of node for the -th facet, where , is a learnable weight matrix of size for each facet, and is the total number of facets. The embeddings for all facets are then concatenated to form a unified multi-facet embedding, , providing a comprehensive representation of the node.
@@ 3.4. Discriminative Facet Consideration using Gumbel Softmax
We utilize the Gumbel Softmax (Jang et al., 2016) function to introduce randomness and uncertainty into facet selection, which helps the model explore diverse facet influences in complex relationships. This approach enables us to capture nuanced facet interactions and dependencies, making our model more adaptable and robust for multi-facet data. Even if the random walk algorithm extracts the paths that are little related to the end nodes, we mitigate their impact by assigning higher weights to the most crucial components using Gumbel Softmax.
In the given equation, represents Gumbel Softmax. is the final embedding for that combines these probabilities, and is a matrix with dimensions .
@@ 3.5. Subgraph Construction
Our approach, however, avoids the need for multiple metapaths by leveraging multi-facet information from a single path between nodes. This results in a subgraph that provides richer information than traditional methods. The MF2Vec model efficiently learns complex node relationships by constructing subgraphs incorporating multi-dimensional edge features which are multi-facet embedding for a more comprehensive node representation. Details are provided in Algorithm 3.
Comparison of HGNN aggregation strategies: HAN, MAGNN, MHNF, MECCH, and MF2Vec for a green target node.
@@ 3.6. Message Passing
To extract node embeddings, we use GCN, considering node interactions and incorporating path facets as edge features. The graph node embedding is computed as follows:
@@ 3.7. Training Losses
We jointly train our model for node classification and link prediction. For node classification, we adopt the cross-entropy loss , where is the set of labeled nodes, is the true label, and is the predicted probability. For link prediction, we use a contrastive loss , where and are the positive and negative node pairs, respectively, and denotes the sigmoid function. These two objectives enable the model to learn both node-level classification and structural link prediction simultaneously.
@@ 4. Experiment
In this part, we assess the performance and practicality of our introduced method through tests on link prediction, node classification, and node clustering tasks across five diverse graph datasets.
Datasets. We evaluate our model on six heterogeneous graph datasets: DBLP, ACM, IMDB, and Freebase for node classification and clustering, and Yelp and MovieLens for link prediction. Dataset details are summarized in Table 2.
@@
Baselines. We compare MF2Vec against SOTA heterogeneous graph neural networks (HGNNs), categorized as follows: (a) Relation-Based HGNNs, including RGCN, RGAT, HGT, HetSANN, and Simple-HGN; and (b) Metapath-Based HGNNs, comprising MP2Vec, HAN, MAGNN, HeCo, MHNF, HGMAE HMSG, and MECCH.
Graph Construction Settings. Following Algorithm 2, we ensured that the node types matched the target node and that the path length was at most 5. For each target node, we performed up to 1000 trials. The number of negative samples was set to 1.
Experimental Settings. To ensure a fair comparison among all HGNNs, the node embedding dimension was fixed to 64. The learning rate was set to , weight decay to , and the dropout rate to 0.5. For each baseline model, hyperparameters such as batch size, attention heads, dropout rate, learning rate, and epochs—excluding the node embedding dimension—followed the default settings specified in the original papers. Training was conducted using the Adam optimizer with early stopping (patience of 20 epochs) to prevent overfitting. Specifically, for MF2Vec, the temperature () was set to 0.5, and the number of facets () was set to 5. To demonstrate generalization, the experiments were conducted using five seed values (1, 10, 100, 1000, and 10000), with the data split into train (80%), validation (10%), and test (10%). The model seeds were fixed during the experiments, and the average and standard deviation of the results were calculated for all seeds.
@@
MF2Vec | 0.9920.002 | 0.9460.011 | 0.9430.010 | 0.9540.008 | 0.8430.008 | 0.8370.008 | 0.7680.033 | 0.6030.041 | 0.5810.052 | 0.8280.023 | 0.6940.021 | 0.6290.047 |
4.1. Downstream Task
4.1.1. Node Classification.
In our node classification experiments, we convert nodes into low-dimensional vector embeddings and apply a Softmax function to obtain probability distributions over label categories. We evaluate the model using Micro-F1, Macro-F1, and AUC metrics, providing a balanced assessment of classification performance. These metrics are used to compare our model, MF2Vec, against baseline models on the DBLP, ACM, IMDB and Freebase datasets, with results presented in Table LABEL:table_1. Our approach consistently outperforms baseline models across all metrics, demonstrating the effectiveness of MF2Vec. The significant performance improvement, particularly on the IMDB dataset, highlights the model’s ability to capture complex relational data through multi-facet vectors. This suggests that MF2Vec effectively reveals nuanced relationships within heterogeneous information networks (HINs), enhancing understanding beyond node types.
4.1.2. Link Prediction.
In Movielens and Yelp, link prediction is framed as a binary classification task to determine the presence of an edge in the original graph. During training, validation, and testing, negative edges are created by replacing the destination node in positive edges with a randomly chosen node that is not connected to the source node. The link probability is computed using the dot product of node embeddings: , where is the sigmoid function. Table 4 presents the AUC, Micro-F1, and Macro-F1 scores for our model and baselines on Yelp and Movielens. MF2Vec performs competitively on Yelp (against MECCH) and outperforms all baselines on MovieLens. These results highlight the advantage of using multi-facet vectors derived from node paths for link prediction, and demonstrate the efficacy of our model in updating embeddings without relying on node types.
@@ 4.1. Downstream Task
4.1.3. Node Clustering.
@@
4.2. Stability under Meta-path Variations
This section empirically evaluates whether MF2Vec can achieve consistent and robust performance across various settings by effectively leveraging multi-facet information shared among all nodes, regardless of the presence or absence of node type-based metapaths. All experiments are conducted on three datasets, and the relative advantages of MF2Vec are assessed through comparison with representative metapath-based GNNs (MAGNN, HMSG, MECCH). The details and interpretations of each experiment are as follows.
4.2.1. Consistency Across Metapath Types
Table 6 presents the performance variability of each model when trained and evaluated on different metapath types (e.g., APA, APCPA, etc.). For each dataset, we measure the standard deviation (Std) of AUC, Mi-F1, and Ma-F1 scores, and find that MF2Vec consistently achieves the lowest standard deviation, indicating the highest stability. This result suggests that MF2Vec does not rely on specific metapaths defined by node types, but instead learns consistent and reliable representations by leveraging shared facet information among nodes.
@@ 4.2.1. Consistency Across Metapath Types
DBLP | AUC | 0.130 | 0.177 | 0.107 | 0.060 |
Mi-F1 | 0.217 | 0.222 | 0.173 | 0.127 | |
Ma-F1 | 0.201 | 0.257 | 0.170 | 0.125 | |
ACM | AUC | 0.038 | 0.039 | 0.034 | 0.032 |
Mi-F1 | 0.078 | 0.058 | 0.051 | 0.033 | |
Ma-F1 | 0.052 | 0.075 | 0.062 | 0.049 | |
IMDB | AUC | 0.034 | 0.039 | 0.041 | 0.025 |
Mi-F1 | 0.020 | 0.023 | 0.049 | 0.008 | |
Ma-F1 | 0.015 | 0.046 | 0.079 | 0.010 |
4.2.2. Generalization Across Unseen Metapaths
Figure 5 compares the generalization performance of each model when trained on a single metapath and tested on different, unseen metapaths. MF2Vec consistently achieves the highest generalization performance even under these unseen conditions, and the performance drop compared to using all metapaths together is significantly smaller than that of other models. This indicates that MF2Vec is less dependent on specific metapath schemas and is able to extract more transferable and robust representations. These results confirm that the ability to select facets independently of explicit path structures is crucial for achieving stable and generalized performance.
Comparison of cross-metapath generalization performance across models. MF2Vec shows smallest performance drop.
Dataset | Yelp | Movielens | DBLP | ACM | IMDB | Freebase | ||||||||||||
Random. | 0.8650.008 | 0.7920.004 | 0.8030.009 | 0.8690.006 | 0.7930.007 | 0.8020.004 | 0.9850.004 | 0.9320.007 | 0.9300.005 | 0.9450.005 | 0.8230.008 | 0.8120.006 | 0.7480.008 | 0.5840.003 | 0.5690.007 | 0.8240.008 | 0.6910.008 | 0.6260.007 |
STE. | 0.8670.018 | 0.7970.003 | 0.8080.015 | 0.8710.006 | 0.7950.006 | 0.8030.005 | 0.9890.003 | 0.9350.009 | 0.9310.01 | 0.9490.009 | 0.8240.013 | 0.8130.012 | 0.7510.005 | 0.5880.004 | 0.5710.003 | 0.8250.020 | 0.6960.017 | 0.6280.042 |
4.3. Time Complexity
@@
The proposed model’s time complexity is , where is the number of nodes, is the average neighbors, is the number of facets, and is the embedding dimension. Attention adds , and graph convolution contributes , resulting in a total complexity of . In Table 8, models like MECCH (Fu and King, 2024), HAN (Wang et al., 2019), and MAGNN (Fu et al., 2020) rely on metapaths, incurring higher costs due to growth with metapath length (). MF2Vec avoids this by leveraging facets (), leading to more efficient computation, satisfying .Additionally, we compare MF2Vec with relation-based methods like HetSANN (Hong et al., 2020) and SimpleHGN (Lv et al., 2021), which also avoid metapaths by focusing on edge types.
Figure 6 compares training times and Macro F1 scores for MF2Vec and other models on DBLP, IMDB, and Movielens. MAGNN (Fu et al., 2020) and MP2Vec (Dong et al., 2017) are excluded from Movielens due to their excessive training times. MF2Vec achieves the best performance on node classification and link prediction tasks, with significantly lower training times. This efficiency arises from its dynamic use of path characteristics, avoiding the high costs of aggregating multiple paths as seen in MAGNN and MP2Vec, ensuring better scalability and faster convergence without compromising performance.
Training time (log seconds) vs Macro F1 across three datasets. MF2Vec achieves best tradeoff with faster convergence.
4.4. Effect of Facet Selection Strategy
Since MF2Vec selects the optimal facet for each path in a learnable manner, we conducted an additional experiment to validate the effectiveness of its facet selection strategy. Table 7 compares various methods for selecting facet information as edge weights. Our GCN-based results show that the proposed method, which leverages Gumbel-Softmax and attention mechanisms, consistently outperforms all alternatives across datasets. When using random facet selection, performance decreases, and omitting facet information entirely leads to the lowest results. We also evaluated the STE (Straight-Through Estimator)-based approach (Bengio et al., 2013), which enables probabilistic gradient propagation during facet selection. However, this method performs slightly worse than our proposed approach. These findings highlight that facet-related information is crucial for effective graph learning, and our method excels at capturing complex structures and hidden relationships.